### v1

@@ VI-B Dynamic Evaluation
Most benchmarks for LLMs safety evaluation remain unchanged after being released by researchers, resulting in mainly static evaluations. This static approach presents three key challenges. First, as LLMs undergo continuous iteration and improvement, many prompts or attack methods that previously triggered unsafe responses or behaviors may now be effectively mitigated and fixed, leading to a "seemingly safe" performance on these benchmarks. Second, data from various safety evaluation datasets may leak, becoming part of LLMs fine-tuning aligned with human values, which could render these datasets ineffective for accurate evaluation. Third, attack methods continue to evolve. For example, jailbreak attacks can disguise malicious prompts that would have been rejected by the model to distract its attention and generate harmful content. As new jailbreak techniques emerge, benchmarks should be updated accordingly to further test model safety mechanisms. Therefore, it is essential to consider dynamic safety evaluation methods in the future, allowing evaluation data to iterate and evolve continuously. Practical implementation can draw on insights from related work on dynamic evaluation for general-purpose LLMs.
@@ VI-C Reliable and Efficient Evaluator
As discussed in Section V, current evaluator approaches fall into three main categories: human-based, rule-based, and model-based. First, while rule-based methods are simple and fast, they have been shown to sometimes mislabel harmful outputs containing predefined strings as harmless, leading to false negatives. Second, although human-based evaluation is considered the gold standard due to its ability to accurately capture linguistic nuances and contextual meanings aligned with real-world values, it is not automated. This method incurs high time and cost demands, making it challenging to scale. Additionally, research has shown that differences in cultural background and moral values among human annotators can lead to significant variability in offense standards[198], affecting the stability and reliability of evaluation results. Furthermore, model-based evaluation offers a new direction for automated evaluation, with advanced models such as GPT-4 showing strong alignment with human judgments and capable of processing large volumes of data efficiently. However, model-based evaluation also faces numerous challenges, including high API costs, significant computational resources needed for model fine-tuning, poor reproducibility of evaluation results, sensitivity to input prompts, and positional biases[213, 206]. These factors can impact the accuracy and fairness of evaluations. Future research should consider how to strike a better balance between model-based and human-based evaluation approaches, achieving reliable evaluation results while maintaining high efficiency and low costs.
VI-D Specific Application Evaluation
@@
As LLMs are increasingly integrated into various real-world applications, evaluating their safety in these specific application tasks has become crucial. For example, [199] developed a physical safety benchmark for evaluating LLMs in drone control systems, demonstrating the importance of targeted safety evaluations within specific domains. However, current safety evaluation efforts for LLMs mainly focus on general models, while safety research for the latest applications remains relatively disconnected, failing to keep pace with the integration of LLMs into other fields. Specifically, many emerging applications such as intelligent medical diagnosis, autonomous driving, weather forecasting, finance, and law have begun leveraging the powerful capabilities of LLMs, yet safety evaluation standards and tools tailored to these contexts are still underdeveloped. This lag not only risks failing to detect and mitigate potential safety hazards in a timely manner but may also limit the further application and development of LLMs in these fields. Therefore, it is urgent to establish safety evaluation frameworks for LLMs in these emerging domains to ensure their safe deployment.
VI-E Multimodal Model And Agent Evaluation
Multimodal large language models (MLLMs) and AI agents, both built upon LLMs, have garnered significant attention in both academia and industry. Multimodal models offer more natural and convenient interactions with humans, while AI agents are used to emulate human-like thinking and decision-making processes. In MLLMs, LLMs play a central role in processing and generating text [201], while in agents, LLMs act as the "brain," handling thought and decision-making. However, the safety evaluation of these two directions currently lags behind their rapid development. Safety evaluation benchmarks for MLLMs remain limited [202, 203, 204, 205], lacking targeted robustness metrics and evaluation standards, and there is insufficient focus on multimodal elements, especially in video-based domains, where evaluations are virtually nonexistent. Additionally, agents introduce numerous new components, creating new attack surfaces and requiring safety assessments in complex interactive environments. While some research has begun to address these challenges [40, 41, 42], there is still no unified consensus on the design standards for the safety benchmarks of the entire AI agent ecosystem [200]. Further safety evaluation issues remain to be solved, such as identifying the core dimensions of agent safety and developing comprehensive methods to evaluate agent performance.
VI-F Beyond Evaluation: Safe to Responsible
@@
Safety evaluation should not be seen as the end goal, but rather as the starting point. It is not only about providing benchmark results for the safe deployment and application of LLMs but also offering direction and insights for future research and development. As the starting point, the core function of safety evaluation is to identify and mitigate potential risks, ensuring the model’s stability and reliability. However, as technology advances and societal demands become more complex, mere safety will no longer suffice to meet high standards for application. Given the current development trends of LLMs and existing research on LLMs safety evaluation, we emphasize that the future evolution of LLMs should transition from merely safe LLMs to more advanced, responsible LLMs. Responsible LLMs should not only possess basic safety, fairness, and trustworthiness but also enhance transparency, interpretability, and auditability. This means that models must not only be technically reliable but also widely recognized for their ethical and social responsibility. Similarly, safety evaluation will evolve and improve alongside advancements in LLMs technology, encompassing a broader range of evaluation dimensions and deeper methods.
@@
By expanding and refining research into the abrove areas, the community can more accurately evaluate the safety of LLMs while also promoting their development towards greater safety. In the future, LLMs will be closely aligned with human societal values and norms, evolving into a responsible form of artificial intelligence.
@@ VII Conclusion
To ensure the safe deployment and responsible utilization of LLMs in real-world applications, safety evaluation is of paramount importance. This paper presents the first systematic and comprehensive review of various aspects of LLMs safety evaluation, structured across four key guiding questions: “Why evaluate”, “What to evaluate”, “Where to evaluate”, and “How to evaluate”. We delve into dimensions of LLMs safety evaluation, including toxicity, robustness, ethics, bias and fairness, and truthfulness, as well as specific downstream tasks, such as privacy, mental health and medical applications, code generation, copyright, and agents. Additionally, we provide current evaluation metrics, toolkits, and an extensive range of datasets and benchmarks for safety evaluation. Furthermore, we summarize the current evaluator roles from three perspectives: human-based, rule-based, and model-based, with the goal of illustrating the state of development in LLMs safety evaluation and offering valuable reference points for researchers in the field. Finally, we highlight several major challenges currently faced in LLMs safety evaluation and discuss potential directions for future research.
This survey is expected to promote the advancement of LLM safety evaluation, providing clear guidance for the safe and controlled use of these models. As safety evaluation techniques continue to progress and improve, we are confident that future LLMs will serve human society more robustly and responsibly. These models will provide strong intelligent support for innovation and development across industries while effectively mitigating potential risks and ensuring a balanced approach between technological progress and trustworthy safety.

### v2

@@ VI-B Dynamic Evaluation
Most benchmarks for LLMs safety evaluation are released as static resources, with limited subsequent updates. This static approach poses several challenges. First, as LLMs evolve through continuous iteration, prompts or attack methods that once exposed vulnerabilities may no longer be effective, resulting in artificially inflated safety performance. Second, evaluation datasets risk contamination, as their contents may be incorporated into model training or fine-tuning for alignment, thereby reducing their validity as independent evaluation resources. Third, adversarial techniques continue to advance; for instance, jailbreak attacks often disguise malicious prompts to bypass safeguards and induce harmful outputs. As new jailbreak strategies emerge, static benchmarks quickly become outdated and insufficient for rigorous safety assessment. Although some initial studies have begun to explore dynamic evaluation [204, 205], this line of research is still nascent. Future efforts should aim to design adaptive safety evaluation methods that enable benchmarks and protocols to evolve in tandem with both model capabilities and adversarial tactics. Promising directions include automated pipelines for generating novel adversarial prompts, community-driven platforms for continuously updating benchmarks, and integrating dynamic evaluation into red-teaming frameworks. Such approaches would help ensure that safety evaluations remain relevant, robust, and reflective of real-world risks.
@@ VI-C Reliable and Efficient Evaluator
As discussed in Section V, current evaluation methods for LLMs safety can be broadly categorized into rule-based, human-based, and model-based approaches. First, rule-based methods are simple and efficient but often fail to capture implicit harms beyond predefined patterns, leading to false negatives. Second, although human-based evaluation is considered the gold standard due to its ability to accurately capture linguistic nuances and contextual meanings aligned with real-world values, it is costly, time-intensive, and difficult to scale. Moreover, cross-cultural and moral value differences among annotators can introduce variability in harm perception and compromise the consistency of results [206]. Third, model-based evaluation offers a new direction for automated evaluation, with advanced models such as GPT-4 showing strong alignment with human judgments and capable of processing large volumes of data efficiently. However, they face significant challenges, including high inference costs, substantial computational requirements for fine-tuning, limited reproducibility, high sensitivity to evaluation prompts, and lack of robustness [197, 199, 200]. These factors can impact the accuracy and fairness of evaluations. Future research should consider how to strike a better balance between model-based and human-based evaluation approaches, achieving reliable evaluation results while maintaining high efficiency and low costs. Future research should prioritize the development of evaluators that are both reliable and efficient. Promising directions include hybrid approaches that combine automated filters with lightweight model-based scoring and selective human oversight, methods to enhance the consistency and reproducibility of model-based evaluators (e.g., prompt calibration and majority vote).
VI-D Application-Specific Evaluation
@@
As LLMs are increasingly integrated into various real-world applications, evaluating their safety in these specific application tasks has become crucial. For example, [207] developed a physical safety benchmark for evaluating LLMs in drone control systems, demonstrating the importance of targeted safety evaluations within specific domains. However, current safety evaluation efforts for LLMs mainly focus on general models, while safety research for the latest applications remains relatively disconnected, failing to keep pace with the integration of LLMs into other fields. General benchmarks, which primarily test for risks such as toxicity or bias, are often insufficient in high-stakes domains, where unique and domain-dependent hazards must be rigorously addressed. Specifically, many emerging applications such as intelligent medical diagnosis, autonomous driving, weather forecasting, finance, and law have begun leveraging the powerful capabilities of LLMs, yet safety evaluation standards and tools tailored to these contexts are still underdeveloped. This gap underscores the urgent need to establish safety evaluation frameworks tailored to emerging domains. Importantly, these frameworks should not only extend beyond generic benchmarks but also incorporate domain expertise and align with regulatory and ethical standards relevant to each field (e.g., FDA regulations in medicine, or financial compliance requirements). Practical implementations could include the creation of domain-specific benchmark datasets, the use of simulation-based environments to test LLM behavior in high-risk scenarios, and the integration of interdisciplinary perspectives to ensure robustness and applicability.
VI-E Safety Evaluation for MLLMs and Agents
Multimodal large language models (MLLMs) and agents, both grounded in LLM technology, have attracted growing attention across academia and industry. MLLMs enable more natural and seamless human–machine interaction by integrating text with modalities such as images, audio, and video, while AI agents emulate human-like reasoning and decision-making processes. In MLLMs, LLMs serve as the core engine for text understanding and generation [208], whereas in agents, LLMs function as the brain, orchestrating planning and decision-making. Despite rapid progress, the safety evaluation of both directions significantly lags behind their development. For MLLMs, existing benchmarks remain limited [209, 210, 211, 212], lacking robustness metrics and evaluation standards tailored to multimodal contexts. In particular, video-based domains such as video understanding and generation remain underexplored. For agents, safety risks are amplified by the integration of multiple new components, which create additional attack surfaces and require evaluation in complex, interactive environments. Examples include emerging communication protocols such as the Model Context Protocol (MCP) and agent-to-agent (A2A) communication, which introduce new possibilities for coordination but also novel vulnerabilities. While some studies have begun addressing these challenges [66, 213, 214], no consensus has yet been reached on unified benchmark design standards for the broader agent ecosystem [215]. Key open challenges include identifying the core dimensions of agent safety, designing evaluation methods capable of capturing risks in dynamic, multi-agent environments, and developing comprehensive frameworks that can evolve alongside these rapidly advancing systems.
VI-F Beyond Evaluation: From Safe to Responsible
@@
Safety evaluation should not be regarded as the ultimate goal, but as the foundation for further progress. It is not only about producing benchmark results to guide the safe deployment of LLMs, but also about shaping directions for future research and development. The core role of safety evaluation is to identify and mitigate risks, thereby ensuring model stability and reliability. Yet, as technology advances and societal demands grow more complex, safety alone will no longer be sufficient to meet the expectations for real-world applications. Building on current development trends and research in LLM safety evaluation, we argue that the future trajectory should move from merely safe LLMs toward truly responsible LLMs. Responsible LLMs should extend beyond basic safety, fairness, and trustworthiness to also embody transparency, interpretability, and auditability. In other words, models must not only be technically reliable, but also broadly acknowledged for their ethical integrity and social responsibility. Correspondingly, safety evaluation itself will continue to evolve, encompassing a wider set of dimensions and employing deeper, more sophisticated methods in order to support this transition.
@@
By expanding and refining research into the above areas, the community can more accurately evaluate the safety of LLMs while also promoting their development towards higher safety standards. In the future, LLMs will be closely aligned with societal values and norms, evolving into a responsible paradigm of artificial intelligence.
@@ VII Conclusion
Ensuring the safe deployment and responsible use of LLMs in real-world applications requires systematic safety evaluation. This paper provides the first comprehensive review of LLMs safety evaluation, organized around four guiding questions: “Why evaluate,” “What to evaluate,” “Where to evaluate,” and “How to evaluate.” We systematically survey key dimensions of LLMs safety, including toxicity, robustness, ethics, bias and fairness, and truthfulness, as well as specific downstream tasks such as privacy, mental health, medical applications, code generation and copyright. In addition, we review current evaluation metrics, and a broad set of datasets and benchmarks, highlighting their characteristics, strengths, and limitations. The roles of different evaluators — human-based, rule-based, and model-based — are also examined to illustrate the current state of LLM safety evaluation. Moreover, we review several representative evaluation frameworks, highlighting how they integrate attacks, defenses, target models, judges in practical assessments. This structured overview aims to provide researchers with a clear understanding of available resources and methodologies, supporting systematic and reproducible safety evaluations. Finally, we highlight several major challenges currently faced in safety evaluation of LLMs and discuss potential directions for future research.
Looking forward, we anticipate that continued research in safety evaluation will facilitate the development of LLMs that are more reliable, responsible, and aligned with human values. By combining robust evaluation frameworks with improved model design, future LLMs can provide intelligent support across diverse applications while effectively mitigating potential risks, ultimately fostering a balanced approach between technological innovation and trustworthy safety.